OpenStudy (anonymous):

dy/dx = 2x - 7, y(2) = 0

OpenStudy (anonymous):

x^2/2 - 7x?

OpenStudy (anonymous):

not sure what to do after that

OpenStudy (rajee_sam):

it is \[\frac{ 2x ^{2} }{ 2 } - 7x + c = x ^{2} - 7x + c\]

OpenStudy (rajee_sam):

Now they have given y(2) = 0. Use this information to find 'c'

OpenStudy (ddcamp):

\[f(x) = 2x - 7, F(x) = x² - 7x + C\] \[F(2) = 0 = 2^2 -7(2) + C = 4 - 14 + C = -10 + C \rightarrow C = 10\] \[F(x) = x^2 - 7x + 10\]

OpenStudy (anonymous):

ty ty

OpenStudy (anonymous):

ok so u did x^2 - 7x + c = 0?

OpenStudy (anonymous):

@rajee_sam